# 3cm3cm2.3cm2.3cm

### v1

@@ 4.2. Examples
For any monomial $f\in{\mathcal{F}}$, we have ${\bf d}_{{\mathcal{F}}}(f)=kf$ for some $k\in{\mathbb{Z}}$, and we define the degree of $f$ to be $k$. In particular, the degree of each ${x^{a}_{i}}$ is $1$, and the degree of each $y^{a}_{r}$ is $-1$. For this reason, it is natural to call ${\bf d}_{{\mathcal{F}}}$ the degree operator on ${\mathcal{F}}$.
@@
${\mathcal{F}}=\bigoplus_{k\in{\mathbb{Z}}}V^{(d)}_{k}.$ (4.9)
@@
$\underline{V\mkern-4.0mu}\mkern 4.0mu{(k_{1},\ldots,k_{{d^{\prime}}})}=V^{(\xi_{1})}_{k_{1}}\otimes\cdots\otimes V^{(\xi_{{d^{\prime}}})}_{k_{{d^{\prime}}}}\qquad\mbox{for \, $k_{1},\ldots,k_{{d^{\prime}}}\in{\mathbb{Z}}$}.$ (4.10)
@@
${\mathcal{F}}=\bigoplus_{k_{1},\ldots,k_{{d^{\prime}}}\in{\mathbb{Z}}}\underline{V\mkern-4.0mu}\mkern 4.0mu{(k_{1},\ldots,k_{{d^{\prime}}})}.$ (4.11)
@@
The matrix
${\rm J}_{k}(\lambda)=\begin{bmatrix}\lambda&1&\ldots&0\\ 0&\lambda&\ddots&\vdots\\ \vdots&\ddots&\ddots&1\\ 0&\ldots&0&\lambda\end{bmatrix}$
is called a $k\times k$ Jordan block with eigenvalue $\lambda$. If $\lambda\not=0$, then
$\big{(}-{\rm J}_{k}(-\lambda)\big{)}^{-1}=\begin{bmatrix}\lambda^{-1}&\lambda^{-2}&\ldots&\lambda^{-k}\\ 0&\lambda^{-1}&\ddots&\vdots\\ \vdots&\ddots&\ddots&\lambda^{-2}\\ 0&\ldots&0&\lambda^{-1}\end{bmatrix}.$ (4.12)
@@
$\mu_{{\boldsymbol{\xi}}}^{{\bf w}}(e_{ab})=-{\rm J}_{\boldsymbol{\xi}}({\bf w})_{a,b}\qquad\mbox{for $a,b=1,\ldots,d$}.$ (4.13)
@@
$\hskip 28.45274pt\nu_{{\boldsymbol{\gamma}}}^{{\bf z}}(E^{i}_{j})=(-1)^{|i|+1}{\rm J}_{{\boldsymbol{\gamma}}}({\bf z})_{i,j}\qquad\mbox{for $i,j\in\mathbb{I}$}.$ (4.14)
@@
$\mathcal{A}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})=\mathcal{A}_{d}^{\mu_{{\boldsymbol{\xi}}}^{{\bf w}}}({\bf z},{\boldsymbol{\gamma}})\quad\text{and}\quad\mathcal{A}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})=\mathcal{A}_{\diamondsuit}^{\nu_{{\boldsymbol{\gamma}}}^{{\bf z}}}({\bf w},{\boldsymbol{\xi}}).$ (4.15)
@@
^$\widehat{{\mathcal{L}}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})=\widehat{{\mathcal{L}}}_{d}^{\mu_{{\boldsymbol{\xi}}}^{{\bf w}}}({\bf z},{\boldsymbol{\gamma}})\quad\text{and}\quad{\mathcal{L}}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})={\mathcal{L}}_{\diamondsuit}^{\nu_{{\boldsymbol{\gamma}}}^{{\bf z}}}({\bf w},{\boldsymbol{\xi}}),$
where the matrices ^$\widehat{{\mathcal{L}}}_{d}^{\mu_{{\boldsymbol{\xi}}}^{{\bf w}}}({\bf z},{\boldsymbol{\gamma}})$ and ${\mathcal{L}}_{\diamondsuit}^{\nu_{{\boldsymbol{\gamma}}}^{{\bf z}}}({\bf w},{\boldsymbol{\xi}})$ are defined as in (3.3) and (3.4).
@@
$\phi:U\big{(}{\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big{)}(\!(\partial_{z}^{-1},z^{-1})\!)\longrightarrow\mathcal{D}(\!(\partial_{z}^{-1},z^{-1})\!),$
@@
respectively.
For $1\leq i\leq{\boldsymbol{\ell}}$, define
$[\gamma_{i}]=\begin{cases}\,\gamma_{i},&\mbox{if}\ \ 1\leq i\leq p^{\prime}\quad\text{or}\quad p^{\prime}+q^{\prime}+1\leq i\leq p^{\prime}+q^{\prime}+m^{\prime};\\ -\gamma_{i},&\mbox{otherwise}.\end{cases}$ (4.16)
We have the following identity.
@@
$\prod_{i=1}^{{\boldsymbol{\ell}}}(\partial_{z}-z_{i})^{[\gamma_{i}]}\cdot\phi\!\left({\rm cdet}\,\!\left(\widehat{{\mathcal{L}}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})\right)\right)=\prod_{a=1}^{{d^{\prime}}}(z-w_{a})^{\xi_{a}}\cdot\varphi\!\left({\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})\right)\right).$
@@
^$\widehat{{\mathcal{L}}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})=J^{t}-\left[\sum_{i=1}^{{\boldsymbol{\ell}}}\sum_{k=0}^{\gamma_{i}-1}\frac{e_{ab}\otimes\overline{t}_{z_{i}}^{k}}{(\partial_{z}-z_{i})^{k+1}}\!\right]^{t}_{\!a,b=1,\ldots,d}$ (4.17)
@@
${\mathcal{L}}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})=J^{\prime}-\left[(-1)^{|i|}\sum_{a=1}^{{d^{\prime}}}\sum_{k=0}^{\xi_{a}-1}\frac{E^{i}_{j}\otimes\overline{t}_{w_{a}}^{k}}{(z-w_{a})^{k+1}}\!\right]_{\!i,j\in\mathbb{I}}.$ (4.18)
@@
$M=\begin{bmatrix}Y^{t}\hskip 5.69054ptP_{X}^{t}\end{bmatrix}\quad\text{and}\quad M^{\prime}=\begin{bmatrix}P_{Y}\vspace{3mm}\\ X\end{bmatrix}.$
@@
$\mathfrak{L}=\begin{bmatrix}J^{t}&M\vspace{2mm}\\ M^{\prime}&J^{\prime}\end{bmatrix}.$
@@
A variant of Corollary 4.9, called the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{\diamondsuit})$, is established in [ChL2, Theorem 4.7].
@@ 4.4. An application of the duality
Every ${\mathfrak{gl}}_{\diamondsuit}({{\bf w},{\boldsymbol{\xi}}})$-module is a ${\mathfrak{gl}}_{\diamondsuit}$-module via the diagonal map from ${\mathfrak{gl}}_{\diamondsuit}$ to ${\mathfrak{gl}}_{\diamondsuit}({{\bf w},{\boldsymbol{\xi}}})$. For any $\mu\in{\mathfrak{h}}_{\diamondsuit}^{*}$ and any ${\mathfrak{gl}}_{\diamondsuit}({{\bf w},{\boldsymbol{\xi}}})$-module $M$, let
$M_{\mu}=\left\{\,\!v\in M\;\middle|\;hv=\mu(h)v\quad\mbox{for all $h\in{\mathfrak{h}}_{\diamondsuit}$}\!\,\right\}.$ (4.19)
@@
${\mathcal{F}}=\bigoplus_{\mu\in{\mathfrak{h}}_{\diamondsuit}^{*}}{\mathcal{F}}_{\mu}.$ (4.20)
Let ${\boldsymbol{\gamma}}=(1^{\boldsymbol{\ell}})$, where ${\boldsymbol{\ell}}=p+q+m+n$. For $1\leq i\leq{\boldsymbol{\ell}}$, define ${\mathfrak{gl}}_{d}^{(i)}={\mathfrak{gl}}_{d}(z_{i},1)\cong{\mathfrak{gl}}_{d}$. Let $W^{(i)}_{\mu_{i}}$ be the subspace of ${\mathbb{C}}[y^{1}_{i},\cdots,y^{d}_{i}]$ (resp., ${\mathbb{C}}[x^{1}_{i},\cdots,x^{d}_{i}]$) spanned by monomials of degree $\mu_{i}\in-{\mathbb{Z}}_{+}$ (resp., ${\mathbb{Z}}_{+}$) if $1\leq i\leq p+q$ (resp., $p+q+1\leq i\leq{\boldsymbol{\ell}}$). We have direct sum decompositions of ${\mathfrak{gl}}_{d}^{(i)}$-modules:
${\mathbb{C}}[y^{1}_{i},\cdots,y^{d}_{i}]=\bigoplus_{\mu_{i}\in-{\mathbb{Z}}_{+}}W^{(i)}_{\mu_{i}}\qquad\hbox{and}\qquad{\mathbb{C}}[x^{1}_{i},\cdots,x^{d}_{i}]=\bigoplus_{\mu_{i}\in{\mathbb{Z}}_{+}}W^{(i)}_{\mu_{i}}.$
@@
${\mathcal{F}}=\bigoplus_{\begin{subarray}{c}\mu_{1},\ldots,\mu_{p+q}\in-{\mathbb{Z}}_{+},\\ \mu_{p+q+1},\ldots,\mu_{\boldsymbol{\ell}}\in{\mathbb{Z}}_{+}\end{subarray}}\underline{W\mkern-4.0mu}\mkern 4.0mu{(\mu_{1},\ldots,\mu_{\boldsymbol{\ell}})},$ (4.21)
@@
$-\sum_{r=1}^{p+q}y^{a}_{r}\partial_{y^{a}_{r}}+\sum_{i=1}^{m+n}(-1)^{|p+q+i|}\partial_{x^{a}_{i}}x^{a}_{i}=-\sum_{r=1}^{p+q}y^{a}_{r}\partial_{y^{a}_{r}}+\sum_{i=1}^{m+n}x^{a}_{i}\partial_{x^{a}_{i}}+(m-n),$
@@
${\mathcal{F}}=\bigoplus_{k_{1},\ldots,k_{{d^{\prime}}}\in{\mathbb{Z}}}{\mathcal{F}}_{[k_{1},\ldots,k_{{d^{\prime}}}]}.$ (4.22)
@@
$\underline{W\mkern-4.0mu}\mkern 4.0mu{(\mu_{1},\ldots,\mu_{\boldsymbol{\ell}})}=\bigoplus_{k_{1},\ldots,k_{{d^{\prime}}}\in{\mathbb{Z}}}\underline{W\mkern-4.0mu}\mkern 4.0mu{(\mu_{1},\ldots,\mu_{\boldsymbol{\ell}})}_{[k_{1},\ldots,k_{{d^{\prime}}}]}.$ (4.23)
@@
[…]
@@
$\mathcal{A}_{\diamondsuit}^{{\bf z},(1^{{\boldsymbol{\ell}}})}({\bf w},{\boldsymbol{\xi}})_{\underline{V\mkern-4.0mu}\mkern 4.0mu{(k_{1},\ldots,k_{{d^{\prime}}})}_{\mu}}=\mathcal{A}_{d}^{{\bf w},{\boldsymbol{\xi}}}({\bf z},(1^{\boldsymbol{\ell}}))_{{\underline{W\mkern-4.0mu}\mkern 4.0mu{(\overline{\mu}_{1},\ldots,\overline{\mu}_{\boldsymbol{\ell}})}}_{[k_{1},\ldots,k_{{d^{\prime}}}]}}.$ (4.24)

### v2

@@ 4.2. Examples
For any monomial $f\in{\mathcal{F}}$, we have ${\bf d}_{{\mathcal{F}}}(f)=kf$ for some $k\in{\mathbb{Z}}$, and we define the degree of $f$ to be $k$. In particular, the degree of each $x^{a}_{i}$} is $1$, and the degree of each $y^{a}_{r}$ is 1-1. For this reason, it is natural to call $\bf d$$_{{\mathcal{F}}}$ the degree operator on $\mathcal{F}$}.
@@
(4.9) ${\mathcal{F}}=\bigoplus_{k\in{\mathbb{Z}}}V^{(d)}_{k}$.
@@
(4.10) $\underline{V\mkern-4.0mu}\mkern 4.0mu{(k_{1},\ldots,k_{{d^{\prime}}})}=V^{(\xi_{1})}_{k_{1}}\otimes\cdots\otimes V^{(\xi_{{d^{\prime}}})}_{k_{{d^{\prime}}}}\qquad\mbox{for \, $k_{1},\ldots,k_{{d^{\prime}}}\in{\mathbb{Z}}$}.$
@@
(4.11) ${\mathcal{F}}=\bigoplus_{k_{1},\ldots,k_{{d^{\prime}}}\in{\mathbb{Z}}}\underline{V\mkern-4.0mu}\mkern 4.0mu{(k_{1},\ldots,k_{{d^{\prime}}})}.$
@@
The $k\times k$ matrix
${\rm J}_{k}(\lambda):=\begin{bmatrix}\lambda&1&\ldots&0\\ 0&\lambda&\ddots&\vdots\\ \vdots&\ddots&\ddots&1\\ 0&\ldots&0&\lambda\end{bmatrix}$
is called a Jordan block with eigenvalue $\lambda$. If $\lambda\not=0$, then
(4.12) $\big(-{\rm J}_{k}(-\lambda)\big)^{-1}=\begin{bmatrix}\lambda^{-1}&\lambda^{-2}&\ldots&\lambda^{-k}\\ 0&\lambda^{-1}&\ddots&\vdots\\ \vdots&\ddots&\ddots&\lambda^{-2}\\ 0&\ldots&0&\lambda^{-1}\end{bmatrix}$.
@@
(4.13) $\mu_{{\boldsymbol{\xi}}}^{{\bf w}}(e_{ab})=-{\rm J}_{\boldsymbol{\xi}}({\bf w})_{a,b}\qquad\mbox{for $a,b=1,\ldots,d$}.$
@@
(4.14) $\hskip 28.45274pt\nu_{{\boldsymbol{\gamma}}}^{{\bf z}}(E^{i}_{j})=(-1)^{|i|+1}{\rm J}_{{\boldsymbol{\gamma}}}({\bf z})_{i,j}\qquad\mbox{for $i,j\in\mathbb{I}$}.$
@@
(4.15) $\mathcal{A}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})=\mathcal{A}_{d}^{\mu_{{\boldsymbol{\xi}}}^{{\bf w}}}({\bf z},{\boldsymbol{\gamma}})\quad\text{and}\quad\mathcal{A}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})=\mathcal{A}_{\diamondsuit}^{\nu_{{\boldsymbol{\gamma}}}^{{\bf z}}}({\bf w},{\boldsymbol{\xi}}).$
@@
$\widehat{{\mathcal{L}}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})=\widehat{{\mathcal{L}}}_{d}^{\mu_{{\boldsymbol{\xi}}}^{{\bf w}}}({\bf z},{\boldsymbol{\gamma}})\quad\text{and}\quad{\mathcal{L}}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})={\mathcal{L}}_{\diamondsuit}^{\nu_{{\boldsymbol{\gamma}}}^{{\bf z}}}({\bf w},{\boldsymbol{\xi}}),$
where the matrices $\widehat{{\mathcal{L}}}_{d}^{\mu_{{\boldsymbol{\xi}}}^{{\bf w}}}({\bf z},{\boldsymbol{\gamma}})$ and ${\mathcal{L}}_{\diamondsuit}^{\nu_{{\boldsymbol{\gamma}}}^{{\bf z}}}({\bf w},{\boldsymbol{\xi}})$ are defined as in (3.3) and (3.4).
@@
$\phi:U\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!),$
@@
respectively. For $1\leq i\leq{\boldsymbol{\ell}}$, define
(4.16) $\gamma_{i}]=\begin{cases}\,\gamma_{i},&\mbox{if}\ \ 1\leq i\leq p^{\prime}\quad\text{or}\quad p^{\prime}+q^{\prime}+1\leq i\leq p^{\prime}+q^{\prime}+m^{\prime};\\ -\gamma_{i},&\mbox{otherwise}.\end{cases}$
@@
We have
$\prod_{i=1}^{{\boldsymbol{\ell}}}(\partial_{z}-z_{i})^{[\gamma_{i}]}\cdot\phi\!\left({\rm cdet}\,\!\left(\widehat{{\mathcal{L}}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})\right)\right)=\prod_{a=1}^{{d^{\prime}}}(z-w_{a})^{\xi_{a}}\cdot\varphi\!\left({\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})\right)\right)$
as an equality between elements of $\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!)$.
@@
(4.17) $\widehat{{\mathcal{L}}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})=J^{t}-\left[\sum_{i=1}^{{\boldsymbol{\ell}}}\sum_{k=0}^{\gamma_{i}-1}\frac{e_{ab}\otimes\overline{t}_{z_{i}}^{k}}{(\partial_{z}-z_{i})^{k+1}}\!\right]^{t}_{\!a,b=1,\ldots,d}$
@@
(4.18) ${\mathcal{L}}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})=J^{\prime}-\left[(-1)^{|i|}\sum_{a=1}^{{d^{\prime}}}\sum_{k=0}^{\xi_{a}-1}\frac{E^{i}_{j}\otimes\overline{t}_{w_{a}}^{k}}{(z-w_{a})^{k+1}}\!\right]_{\!i,j\in\mathbb{I}}.$
@@
$M=\begin{bmatrix}Y^{t}\hskip 5.69054ptP_{X}^{t}\end{bmatrix}\quad\text{and}\quad M^{\prime}=\begin{bmatrix}P_{Y}\vskip 8.53581pt\\ X\end{bmatrix}$.
@@
$\mathfrak{L}=\begin{bmatrix}J^{t}&M\vskip 5.69054pt\\ M^{\prime}&J^{\prime}\end{bmatrix}$.
@@
A variant of Corollary 4.9, called the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{\diamondsuit})$, is established in [ChL2].
@@ 4.4. An application of the duality
[…]
(4.19) $M_{\mu}=\left\{\,\!v\in M\;\middle|\;hv=\mu(h)v\quad\mbox{for all $h\in{\mathfrak{h}}_{\diamondsuit}$}\!\,\right\}.$
@@
(4.20) ${\mathcal{F}}=\bigoplus_{\mu\in{\mathfrak{h}}_{\diamondsuit}^{*}}{\mathcal{F}}_{\mu}.$
Let ${\boldsymbol{\gamma}}=(1^{\boldsymbol{\ell}})$, where ${\boldsymbol{\ell}}=p+q+m+n$. For $1\leq i\leq{\boldsymbol{\ell}}$, define ${\mathfrak{gl}}_{d}^{(i)}={\mathfrak{gl}}_{d}(z_{i},1)\cong{\mathfrak{gl}}_{d}$. Let $W^{(i)}_{\mu_{i}}$ be the subspace of ${\mathbb{C}}[y^{1}_{i},\cdots,y^{d}_{i}]$ (resp., ${\mathbb{C}}[x^{1}_{i-p-q},\cdots,x^{d}_{i-p-q}]$) spanned by monomials of degree $\mu_{i}\in-{\mathbb{Z}}_{+}$ (resp., $\mathbb{Z}$$_{+}$) if $1\leq i\leq p+q$ (resp., $p+q+1\leq i\leq{\boldsymbol{\ell}}$). We have direct sum decompositions of ${\mathfrak{gl}}_{d}^{(i)}$-modules:
${\mathbb{C}}[y^{1}_{i},\cdots,y^{d}_{i}]=\bigoplus_{\mu_{i}\in-{\mathbb{Z}}_{+}}W^{(i)}_{\mu_{i}},\qquad 1\leq i\leq p+q,$
and
${\mathbb{C}}[x^{1}_{i-p-q},\cdots,x^{d}_{i-p-q}]=\bigoplus_{\mu_{i}\in{\mathbb{Z}}_{+}}W^{(i)}_{\mu_{i}},\qquad p+q+1\leq i\leq{\boldsymbol{\ell}}.$
@@
(4.21) ${\mathcal{F}}=\bigoplus_{\begin{subarray}{c}\mu_{1},\ldots,\mu_{p+q}\in-{\mathbb{Z}}_{+},\\ \mu_{p+q+1},\ldots,\mu_{\boldsymbol{\ell}}\in{\mathbb{Z}}_{+}\end{subarray}}\underline{W\mkern-4.0mu}\mkern 4.0mu{(\mu_{1},\ldots,\mu_{\boldsymbol{\ell}})},$
@@
$\sum_{r=1}^{p+q}y^{a}_{r}\partial_{y^{a}_{r}}+\sum_{i=1}^{m+n}(-1)^{|p+q+i|}\partial_{x^{a}_{i}}x^{a}_{i}=-\sum_{r=1}^{p+q}y^{a}_{r}\partial_{y^{a}_{r}}+\sum_{i=1}^{m+n}x^{a}_{i}\partial_{x^{a}_{i}}+(m-n),$
@@
(4.22) ${\mathcal{F}}=\bigoplus_{k_{1},\ldots,k_{{d^{\prime}}}\in{\mathbb{Z}}}{\mathcal{F}}_{[k_{1},\ldots,k_{{d^{\prime}}}]}.$
@@
(4.23) $\underline{W\mkern-4.0mu}\mkern 4.0mu{(\mu_{1},\ldots,\mu_{\boldsymbol{\ell}})}=\bigoplus_{k_{1},\ldots,k_{{d^{\prime}}}\in{\mathbb{Z}}}\underline{W\mkern-4.0mu}\mkern 4.0mu{(\mu_{1},\ldots,\mu_{\boldsymbol{\ell}})}_{[k_{1},\ldots,k_{{d^{\prime}}}]}.$
@@
${\mathcal{F}}_{[k_{1},\ldots,k_{d^{\prime}}]}=\underline{V\mkern-4.0mu}\mkern 4.0mu{(k_{1},\ldots,k_{{d^{\prime}}})}.$
@@
(4.24) $\mathcal{A}_{\diamondsuit}^{{\bf z},(1^{{\boldsymbol{\ell}}})}({\bf w},{\boldsymbol{\xi}})_{\underline{V\mkern-4.0mu}\mkern 4.0mu{(k_{1},\ldots,k_{{d^{\prime}}})}_{\mu}}=\mathcal{A}_{d}^{{\bf w},{\boldsymbol{\xi}}}({\bf z},(1^{\boldsymbol{\ell}}))_{{\underline{W\mkern-4.0mu}\mkern 4.0mu{(\overline{\mu}_{1},\ldots,\overline{\mu}_{\boldsymbol{\ell}})}}_{[k_{1},\ldots,k_{{d^{\prime}}}]}}.$